# A2ATS: Retrieval-Based KV Cache Reduction via Windowed Rotary Position Embedding and Query-Aware Vector Quantization

### v1

@@ 1 Introduction
Various methods were proposed to address these challenges from different perspectives. Quantization-based methods (Liu et al., 2024b; Hooper et al., 2024) compress KV cache by using lower bit-width representations for KV cache elements. Eviction-based methods (Xiao et al., 2024; Zhang et al., 2023; Li et al., 2024; Yang et al., 2024b) reduce the KV cache size by directly evicting unimportant tokens from memory. Retrieval-based methods (Tang et al., 2024; Singhania et al., 2024; Zhang et al., 2024; Liu et al., 2024a; Chen et al., 2024) offload the complete KV cache to CPU memory and retrieve necessary tokens on demand during inference. However, these methods still face challenges of limited compression ratio, unsatisfactory accuracy degradation, or extra retrieval overhead.
@@
Experimental results demonstrate that A2ATS can achieve a low accuracy degradation of 2.2 on Llama-3.1-8B and 0.4 on Mistral-7B while accessing only 6% of the entire KV cache, thereby increasing long context serving throughput by up to 2.7×2.7$\times$2.7 ×. Our source code is publicly available  [^1] https://anonymous.4open.science/r/4987d19d-f5cd-4f14-910d-c7141ce37f13/.
@@ 3.1 Inter-Input Similarity of Codebooks
We utilize the Llama-3.1-8B-Instruct model (Dubey et al., 2024), and two random samples from the FineWeb dataset (Penedo et al., 2024) with a context length of approximately 32k tokens each for experiments. We collect pre-PE and post-PE (RoPE is used in here) key states from all attention heads on both samples, then employ k-means++ algorithm (Arthur and Vassilvitskii, 2007) to generate codebooks with a size of 4096 codewords for each set of key states, and calculate the inter-sample cosine similarities of pre-PE and post-PE codebooks using Equation (7). The similarity generated by pre-PE and post-PE codebooks are shown in Figure 1. From the experimental results, we derive the following key observations:
@@ 5.1 Experimental Setup
Methods. For main experiments, we compare the proposed A2ATS with the following four KV cache reduction methods, along with the full attention baseline: H2O (Zhang et al., 2023), SnapKV (Li et al., 2024), Quest (Tang et al., 2024), MagicPIG (Chen et al., 2024). For a fair comparison, the sparsity ratios of all KV cache reduction methods are controlled around 0.06. Detailed discussions are presented in Appendix A.
@@ 5.3 Ablation Study
Table 2 validates the effectiveness of WRoPE and query-aware vector quantization on improving model accuracy. Experimental results draw the following conclusions: (1) WRoPE is fundamental to attention score approximation using shared codebooks. (2) Query-aware vector quantization provides a further improvement in model accuracy by aligning the objectives of vector quantization and attention score approximation. Detailed discussions are presented in Appendix B.
@@ 6 Related Work
Retrieval-based KV cache reduction. This method keeps the entire KV cache in memory while selectively retrieving tokens crucial for the current inference. Quest (Tang et al., 2024) chunks the continuous KV cache into pages and pre-calculates necessary metadata for each page during prefilling. For decoding, it selects the top-K critical cache pages to participate in selective attention computation. PQCache (Zhang et al., 2024) and ClusterKV (Liu et al., 2024a) perform vector quantization on the key states with individual codebooks constructed for each input during prefilling. For decoding, the system uses codewords of the codebooks to approximate attention scores, then retrieves the top-K tokens for computation. MagicPIG (Chen et al., 2024) employs Locality-Sensitive Hashing on query and key states for token retrieval. Although these methods generally achieve lower performance degradation compared to eviction-based methods, they still suffer from unsatisfactory performance and lead to extra overhead.
@@ 7 Conclusion
In this paper, we propose A2ATS, a novel retrieval-based KV cache reduction method. First, we propose Windowed Rotary Position Embedding to decouple the positional dependency from query and key states after position embedding. Then, we propose query-aware vector quantization to achieve an accurate attention score approximation. Next, we introduce the heterogeneous inference design for KV cache offloading which increases available batch size. Experimental results demonstrate that A2ATS achieves lower performance degradation with comparable or lower overhead compared to existing methods, thereby boosting long context serving throughput by up to 2.7×2.7$\times$2.7 ×.
@@ Appendix A Method Configurations
For a fair comparison, the sparsity ratios of all KV cache reduction methods are controlled around 0.06, which means approximately 6% of KV cache is accessed at each inference. It is important to note that the sparsity ratio discussed in this paper differs in definition from the $\mathrm{cost}_{2}$ in MagicPIG (Chen et al., 2024). Specifically, $\mathrm{cost}_{2}$ measures the ratio of computation overhead (FLOPs) compared to full attention, whereas our sparsity ratio measures the ratio of memory access overhead (MOPs) relative to full attention. Since attention modules are typically considered memory-bound (Kim et al., 2023), we argue that the latter metric provides more meaningful insights on potential overhead reduction. Additionally, the initial 4 tokens and the most recent 64 tokens are statically preserved to align with MagicPIG (Xiao et al., 2024). The detailed configurations for each methods are shown in Table 3.
@@
Appendix B Detailed Discussions of Ablation Study
@@
Query-aware vector quantization provides a further improvement in model accuracy by aligning the objectives of vector quantization and attention score approximation. Our full method, incorporating query-aware vector quantization, demonstrates further improvements, particularly at longer context lengths (+1.5 at 64K, +1.2 at 96K, respectively). However, query-aware vector quantization alone underperforms the baseline, exhibiting a more significant drop at longer context lengths. This performance degradation suggests that representation divergence is not solely limited to key states but also affects query states, hindering the effectiveness of query-aware vector quantization. With WRoPE mitigating positional dependencies, query-aware vector quantization further optimizes the attention score approximation, achieving state-of-the-art performance.

### v2

@@ Abstract
footnotetext: *Equal contributions.footnotetext: Corresponding authors.
@@ 1 Introduction
Various methods were proposed to address these challenges from different perspectives. Quantization-based methods (Liu et al., 2024b; Hooper et al., 2024) compress KV cache by using lower bit-width representations for KV cache elements. Eviction-based methods (Xiao et al., 2024; Zhang et al., 2023; Li et al., 2024; Yang et al., 2024b) reduce the KV cache size by directly evicting unimportant tokens from memory. Retrieval-based methods (Tang et al., 2024; Singhania et al., 2024; Zhang et al., 2024; Liu et al., 2024a; Chen et al., 2025) offload the complete KV cache to CPU memory and retrieve necessary tokens on demand during inference. However, these methods still face challenges of limited compression ratio, unsatisfactory accuracy degradation, or extra retrieval overhead.
@@
Experimental results demonstrate that A2ATS can achieve a low accuracy degradation of 2.2 on Llama-3.1-8B and 0.4 on Mistral-7B while accessing only 6% of the entire KV cache, thereby increasing long context serving throughput by up to 2.7×2.7$\times$2.7 ×. Our source code is publicly available1.
1https://github.com/junhuihe-hjh/A2ATS
@@ 3.1 Inter-Input Similarity of Codebooks
We utilize the Llama-3.1-8B-Instruct model (Dubey et al., 2024), and two random samples from the FineWeb dataset (Penedo et al., 2024) with a context length of approximately 32k tokens each for experiments. We collect pre-PE and post-PE (RoPE is used in here) key states from all attention heads on both samples, then employ k-means++ algorithm (Arthur and Vassilvitskii, 2007) to generate codebooks with a size of 4096 codewords for each set of key states, and calculate the inter-sample cosine similarities of pre-PE and post-PE codebooks using Equation 7. The similarities obtained are shown in Figure 1. From the experimental results, we derive the following key observations:
@@ 5.1 Experimental Setup
Methods. For main experiments, we compare the proposed A2ATS with the following four KV cache reduction methods, along with the full attention baseline: H2O (Zhang et al., 2023), SnapKV (Li et al., 2024), Quest (Tang et al., 2024), MagicPIG (Chen et al., 2025). For a fair comparison, the sparsity ratios of all KV cache reduction methods are controlled around 0.06. Detailed discussions are presented in Appendix A.
@@ 5.3 Ablation Study
Table 2 validates the effectiveness of WRoPE and query-aware vector quantization on improving model accuracy. Experimental results draw the following conclusions: (1) WRoPE is fundamental to attention score approximation using shared codebooks. (2) Query-aware vector quantization provides a further improvement in model accuracy by aligning the objectives of vector quantization and attention score approximation. Detailed analysis is presented in Appendix B.
@@ 6 Related Work
Retrieval-based KV cache reduction. This method keeps the entire KV cache in memory while selectively retrieving tokens crucial for the current inference. Quest (Tang et al., 2024) chunks the continuous KV cache into pages and pre-calculates necessary metadata for each page during prefilling. For decoding, it selects the top-K critical cache pages to participate in selective attention computation. PQCache (Zhang et al., 2024) and ClusterKV (Liu et al., 2024a) perform vector quantization on the key states with individual codebooks constructed for each input during prefilling. For decoding, the system uses codewords of the codebooks to approximate attention scores, then retrieves the top-K tokens for computation. MagicPIG (Chen et al., 2025) employs Locality-Sensitive Hashing on query and key states for token retrieval. Although these methods generally achieve lower performance degradation compared to eviction-based methods, they still suffer from unsatisfactory performance and lead to extra overhead.
@@ 7 Conclusion
In this paper, we propose A2ATS, a novel retrieval-based KV cache reduction method. First, we propose Windowed Rotary Position Embedding to decouple the positional dependency from query and key states after position embedding. Then, we propose query-aware vector quantization to achieve an accurate attention score approximation. Next, we introduce the heterogeneous inference design for KV cache offloading, which increases available batch size. Experimental results demonstrate that A2ATS achieves lower performance degradation with comparable or lower overhead compared to existing methods, thereby boosting long context serving throughput by up to 2.7×2.7$\times$2.7 ×.
@@ 8 Limitations
Acknowledgements
We thank all the reviewers for their insightful comments. This work is supported by the National Natural Science Foundation of China (No. 62472330).
@@ Appendix A Method Configurations
For a fair comparison, the sparsity ratios of all KV cache reduction methods are controlled around 0.06, which means approximately 6% of KV cache is accessed at each inference. It is important to note that the sparsity ratio discussed in this paper differs in definition from the $\mathrm{cost}_{2}$ in MagicPIG (Chen et al., 2025). Specifically, $\mathrm{cost}_{2}$ measures the ratio of computation overhead (FLOPs) compared to full attention, whereas our sparsity ratio measures the ratio of memory access overhead (MOPs) relative to full attention. Since attention modules are typically considered memory-bound (Kim et al., 2023), we argue that the latter metric provides more meaningful insights on potential overhead reduction. Additionally, the initial 4 tokens and the most recent 64 tokens are statically preserved to align with MagicPIG (Xiao et al., 2024). The detailed configurations for each methods are shown in Table 3.
@@
Appendix B Detailed Ablation Study Analysis
@@
Query-aware vector quantization provides a further improvement in model accuracy by aligning the objectives of vector quantization and attention score approximation. Our full method, incorporating query-aware vector quantization, demonstrates further improvements, particularly at longer context lengths (+1.5 at 64K, +1.2 at 96K, respectively). However, query-aware vector quantization alone does not outperform conventional vector quantization. This is because query-aware vector quantization relies on the estimation of the second-moment matrix of queries, i.e., $H=\mathbb{E}_{\tilde{q}\sim\mathcal{D}^{\text{query}}}[\tilde{q}^{\top}\tilde{q}]$ (Section 4.2, Equation 13). Positional dependencies induced by RoPE cause representation divergence in $\tilde{q}^{\top}\tilde{q}$, hindering the accurate estimation of $H$ and leading to decreased accuracy. The synergy between WRoPE (which eliminates positional dependencies) and query-aware quantization (which aligns the objective between vector quantization and attention score approximation) ultimately delivers state-of-the-art performance.
Appendix C Additional Experiments on LongBench
To further validate the effectiveness of our proposed method on real-world tasks, we conducted additional experiments on 6 tasks from LongBench (Bai et al., 2024) (HotpotQA, MultiFieldQA-en, QMSum, TriviaQA, PassageRetrieval-en and RepoBench-P). The configurations of all KV cache reduction methods are the same as those presented Section 5.1.
[…]
HpQA MfQA QMS TrQA PRe RBP Average
Llama-3.1-8B-Instruct 1.00 58.41 56.43 25.01 91.47 100.0 56.37 64.62
H2O 0.060 57.44 49.72 24.09 91.55 99.50 52.28 62.43
SnapKV 0.060 57.38 54.12 24.59 90.92 99.50 54.36 63.48
Quest 0.060 57.79 55.49 24.58 90.70 99.50 53.94 63.67
MagicPIG 0.068 57.80 55.99 25.26 90.82 99.50 55.31 64.11
A2ATS 0.060 58.03 56.84 25.03 91.63 99.50 56.27 64.55
Table 4: Experimental results on LongBench.
Table 4 compares the accuracies of different methods, along with attention sparsity ratios. Experimental results demonstrate that the proposed A2ATS outperforms other baselines under comparable sparsity, emphasizing the effectiveness of our proposed method on a broader range of long-context tasks.